# Banyan: Improved Representation Learning with Explicit Structure

### v3

@@ Abstract
We present Banyan, a model that efficiently learns semantic representations by leveraging explicit hierarchical structure. While transformers excel at scale, they struggle in low-resource settings. Conversely recent structured models have shown promise as efficient learners, but lack performance. Banyan bridges this gap with two key innovations: an entangled hierarchical tree structure and diagonalized message passing, enabling it to outperform larger transformer models with just 14 non-embedding parameters. It excels in low-resource settings, offering a viable alternative for under-represented languages and highlighting its potential for efficient, interpretable NLP in resource-constrained environments.
@@ 1 Introduction
Banyan, achieves performance comparable to transformer-based baselines and represents a low-cost, viable alternative to transformers for producing representations in low-resource languages, as measured by semantic textual similarity (STS) tasks. By leveraging cognitively inspired inductive biases, our work enables semantic representation learning that rivals or surpasses large-scale pre-trained LLMs—using only 14 non-embedding parameters. This breakthrough offers a new direction for efficient and effective semantic understanding in resource-constrained environments.
@@ 2 Background and Related Work
Recursive Neural Networks (RvNNs): Like regular recurrent neural networks (RNNs), RvNNs process data by repeatedly applying a function to update their state in sequence. However, instead of relying on temporal ordering (like the sequence of words in a sentence), RvNNs use hierarchical structures, often provided as input—most commonly as a binary tree—and can be applied either bottom-up (from leaves to root) or top-down (root to leaves). They were first introduced by Socher et al. (2011, 2013), inspiring many successor models that vary in how they define the recursive function, including Tree-LSTMs (Tai et al., 2015) and IORNN (Le & Zuidema, 2014).
@@
Semantic Representations of Text: Systems like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) use the distributional hypothesis (Harris, 1954) to model word semantics, which posits that words are defined by the context in which they appear. To learn representations, these models use a fixed context window and predict a missing word in a sequence. While initially effective, this approach has limitations because words can have multiple meanings depending on context. Early encoder-only transformer models produced poor representations (Reimers & Gurevych, 2019), but recent developments with contrastive finetuning (Gao et al., 2021) have become the method of choice for producing semantic representations.
Semantic Representation Learning through Structure: Transformer embeddings have become more successful than static word embeddings due to their ability to handle varying contextual influences. Unlike attention mechanisms in transformers, which route information based on token relationships, some approaches use explicit graphs or structures—such as dependency (Levy & Goldberg, 2014; Vashishth et al., 2019) or constituency (Opper et al., 2023b) parses—to determine the focus of context windows. Opper et al. (2023b) uses constituency parsers to learn sentence-level embeddings alongside word embeddings, introducing two models: StrAE, which relies on being given its structure, and Self-StrAE, which learns its own structure using representations. The latter model, Self-StrAE, serves as the foundation for Banyan and is described next.
@@ 4 Banyan
Figure 2: Entangled trees: Example of disjoint trees being transformed into an entangled tree. Internal functions ($C_{\Phi},D_{\Theta},\ldots$) are elided to avoid clutter.
@@ 4.1 From trees to entangled trees
We wish to have composition embeddings amortise over all possible contexts, and simultaneously, all decompositions embeddings to resolve to the same thing. The representation of an entity Lisa should encapsulate everything she could possibly eat. Simultaneously, the average of everything she could eat we should get back to Lisa. Self-StrAE does not explicitly model this behaviour in its structure. Decomposition embeddings of the same entity only interact when we calculate the loss. On top of this, because the loss is taken over the batch, they are actually treated as false negatives to each other. Even though they are terms that ought not be pushed away, the objective ask them to be.
@@ 4.2 Simplified Message Passing
with sigmoid non-linearity ($\sigma$) applied to parameters both for numerical stability and to enforce a decayed memory over structure depth. Repeated application of the diagonal composition function will decay the influence of nodes further down in the tree, thereby respecting the compression order of the structure. Also, during composition parent representations can increase in magnitude as they are the sum of the two children. During decomposition child representations will, by necessity, reduce back down in magnitude towards the leaves. In this way the functions further mimic the information flow specified by the entangled trees.
These relatively simple changes have a pretty drastic effect, both in terms of performance (see experiments) as well as memory footprint, with parameters now reduced by a factor of $U$ compared to the functions from Equations 1 and 2
Table 1: Sentence level results for models pretrained on English. Higher is better. Results averaged across four random initialisations. Bolded results indicate no standard deviation overlap. Spearman’s $\rho$ is * 100 following convention.
@@
Banyan 16.57 $\pm$ 0.02 63.25 $\pm$ 0.03 69.00 $\pm$ 0.01 49.61 $\pm$ 0.02
5 Experiments:
5.1 Warmup: English language evaluation
Goal: We wish to test whether Banyan can efficiently learn semantics. We start by evaluating on English, as there are far more test sets available than for low resource languages.
Evaluation: We want to evaluate how well Banyan learns effective semantic representations. Ideally we want to probe this at different levels of hierarchy, covering both the lexical and sentential level. Our evaluation is unsupervised, both to directly probe the effect of pretraining with the inductive bias, and because this setting has greater parity to what may be expected in a low-resource domain, with few labelled datasets. For these reasons, we turn to a series of tasks which measure correlation between cosine similarity of embedding pairs for two examples and human judgements of their semantic correspondence. On the word level, we use Simlex-999 (Hill et al., 2015) and WordSim-S/R (Agirre et al., 2009). All tasks measure semantics, but do so on differing axes. Simlex measures similarity at the exclusion of relatedness. Wordsim S measures similarity without penalising relatedness. And Wordsim R measures relatedness. On the sentence level, we use STS-12 through 16 (Agirre et al., 2012, 2013, 2014, 2015, 2016), the STS-B (Cer et al., 2017), SICK-R (Marelli et al., 2014) and SemRel (Ousidhoum et al., 2024). In addition to semantic similarity we look at retrieval and classification tasks to ensure embedding robustness. We use the Quora and Arguana retrieval datasets from the BEIR suite (Thakur et al., 2021) alongside MRPC and SST-2 from the GLUE benchmark. For the classification tasks we train a GeLU MLP with a 512D hidden size, and leave models frozen to emphasise the effect of pre-training. A good embedding model should do well across the board.
@@ 5.1 Warmup: English language evaluation
Baselines: We compare against the Self-StrAE, GloVe embeddings (Pennington et al., 2014) and a RoBERTa (Liu et al., 2019) in the medium configuration from (Turc et al., 2019; Opper et al., 2023a). Self-StrAE stands as the closest point of comparison to Banyan. Self-StrAE indicates the performance level of structured representation learning lies, as well as any improvements we are able to achieve. GloVe lets us compare to traditional static embeddings. This comparison probes whether our model is learning anything more than just simple bag of word features. To obtain sentence embeddings, we report results using both the simple average of the word embeddings and the average with stopwords removed following (Reimers & Gurevych, 2019). The latter is generally stronger. For RoBERTa, we report results using both the standard model, and again after enhancing RoBERTa through an extra round of contrastive SimCSE training (Gao et al., 2021), as a further STS baseline. Our pooling strategy is mean. To produce static embeddings from RoBERTa to use in lexical evaluation, we follow Bommasani et al. (2020) and average the contextualised representations of all occurrences of the word in the training set. The RoBERTa is intended as a stronger baseline. It has significantly more parameters than Banyan and is able to model meaning in context unlike GloVe.
@@
Hyperparameters and Pre-training Details: For all models we set the embedding size to 256. For Self-StrAE we use the configuration of (Opper et al., 2023b) and set embeddings as square matrices (i.e., $K$=16 and $U$=16). For Banyan we set these values to $K$=128 and $U$=2, because the more independent channels we allowed the better the model seemed to perform. We refer the reader to the Appendix for ablations. We also note that because we can perform this reduction in channel size, the number of non-embedding parameters for Banyan drops to just 14, as these are directly proportional to $U$. We trained Self-StrAE and Banyan for 15 epochs (circa 15k steps and sufficient for convergence) using the Adam optimizer (Kingma & Ba, 2015), with a learning rate of 1e-3 for Banyan and 1e-4 for Self-StrAE  using a batch size of 512. To process the graphs we used DGL (Wang et al., 2020). The GloVe baseline was trained for 15 epochs with a learning rate of 1e-3, and a window size of 10. We used the official C++ implementation. RoBERTa medium was trained for 200,000 steps, (10% of which were used for warmup). We used a learning rate of 5e-5, and a linear schedule. Positional embeddings are relative key-query. The configuration for RoBERTa medium is 8 layers, 8 attention heads and 2048 dimensional feedforward layers. We used the Transformers library to implement and train the model (Wolf et al., 2020). For SimCSE training, we used the default parameters and the official implementation for unsupervised RoBERTa training from Gao et al. (2021). As our pre-training corpus we selected the WikiText-103 benchmark dataset (Merity et al., 2016). The RoBERTa and GloVe baselines are trained on the full corpus (103 million tokens), representing the upper-middle end of the level of data scale that might be available for a language. Whereas we trained Self-StrAE and Banyan on a uniform subsample of 10 million tokens, representing the lower end of how many tokens might be available, because these explicit structure models are supposed to be efficient learners.
@@
Results: Results are shown in Tables 1, 2 and 3. On both the word level and sentence level Banyan does much better than Self-StrAE. We ablate the reasons for this in more detail later in the manuscript. Both models suffer on SimLex because they need to model both similarity and relatedness as the latter dictates merge (related concepts often compose together). However, the important thing to note is that the structured models effectively transfer the same performance from the word level to the sentence level. They can take advantage of composition, and transfer the meaning of the parts to understanding the meaning of the whole. The GloVe baseline is good on the word level, but does not generalise to the sentence level as well as the transformer, even when we give it stopword removal. It cannot transfer semantic knowledge seamlessly to different levels of complexity. Banyan can, and is able to achieve comparable or better performance than the SimCSE RoBERTa despite being much smaller and exposed to 10x less pre-training data. This means we have a structured model that remains efficient and cheap, and also effective at representation learning.
@@
Model Indonesian Arabic Telugu Marathi Mor. Arabic Kinyarwanda Hausa Afrikaans Spanish Amharic Hindi Score
XLM-R 46.7 31.6 46.3 55.7 17.4 13.2 4.1 56.2 68.9 57.3 52.7 40.92
Llama-3.1 (8B) 53.4 31.1 65.6 63.4 19.4 19.7 6.1 65.4 66.7 64.1 61.7 46.96
Mistral Nemo 50.7 20.1 57 52.3 15.1 16.3 1.8 58.3 66.2 53.2 55.8 40.62
@@
XLM-R (FT) 47.9 33.6 68.8 75.1 21.6 19.4 14.6 72.6 72.8 59.6 57.6 49.41
Banyan 44.17 43.20 71.13 67.67 52.00 46.1 43.7 78.68 60.95 66.18 61.83 57.78
$\pm$ 1.11 $\pm$ 1.82 $\pm$ 0.91 $\pm$ 0.64 $\pm$ 2.25 $\pm$ 0.32 $\pm$ 1.21 $\pm$ 0.30 $\pm$ 0.76 $\pm$ 0.46 $\pm$ 0.6
5.2 Multilingual Evaluation:
@@
Goal: From the results in English we know that Banyan is an efficient learner: it can produce good representations without requiring large-scale data or compute. This implies potential use for under represented communities, whose languages are not well covered by current NLP approaches. Now we have to test that.
@@
Evaluation: Learning semantic representations for low-resource languages remains an ongoing challenge in NLP. A core problem is not just the lack of training data, but also the lack of evaluation datasets. Recent work by Ousidhoum et al. (2024) has sought to address this issue, providing semantic relatedness test sets for several low resource Asian and African languages. These test sets are evaluated the same as before, comparing the cosine similarity between model embeddings for two sequences with human judgements of their semantic match. As before, evaluation is zero-shot unsupervised, allowing us to evaluate Banyan on Indonesian, Arabic, Telugu, Marathi, Moroccan Arabic, Kinyarwanda, Hausa, Afrikaans, Spanish, Amharic and Hindi—covering a broad spectrum of resource extent. For example, Spanish and Hindi are reasonably well represented, while Moroccan Arabic and Kinyarwanda have extremely little training data.
@@
Baselines: We select XLM-R (Conneau et al., 2019): a transformer encoder trained on 2TB of multilingual data. Llama 3.1 8B (Dubey et al., 2024): a decoder only LLM trained on 15 trillion tokens. Mistral Nemo 12B: a decoder only LLM designed with multi-lingual capacities in mind. In addition we also compare against two specialised embedding models from the sentence transformers range (Reimers & Gurevych, 2019): Mini-LM-L12-V2 and Paraphrase-XLM-R-Multilingual-V1. These are pre-trained transformer encoders that have been finetuned on supervised datasets designed to produce high quality semantic representations. The baselines we select here are emblematic of the kind of models one might reach for in order to embed a corpus. For all models we use mean pooling to produce the sentence representation following (Reimers & Gurevych, 2019; Li & Li, 2024). Finally, for parity we include an XLM-R baseline which is finetuned on the same corpora.
@@
Banyan Pre-training and Hyperparameters: For Afrikaans, Spanish and Amharic we obtained corpora from Leipzig Corpora Collection [^1] For Spanish and Hindi we select the mixed corpus and uniformly subsample to reduce size to $\approx$10M tokens. (Goldhahn et al., 2012). For Amharic we utilised a MiT licenced pre-training set of 1 million sequences available on the Huggingface hub at this link. Kinyarwanda and Hausa data was sourced from Opus (Nygaard & Tiedemann, 2003). Each dataset consists of roughly 10 million tokens. We utilise a pre-trained BPE tokenizer for each language from the BPEMB Python package (Heinzerling & Strube, 2018). For the model hyperparameters we keep all the settings from the experiments on English. For XLM-R we finetune for up to 100k steps with early stopping, using a linearly scheduled learning rate of 5e-5 with 10% of steps as warmup. XLM-R runs at batch size 128 across 4xA40 45gb cards.
Results: See Table 4. In Spanish, a well-resourced language with high coverage, the transformer baselines almost all outperform Banyan. However, as languages become lower resourced the picture changes, and Banyan outperforms or is comparable to the baselines. This even includes the multilingual XLM-R that has undergone supervised training. While finetuning XLM-R improves performance the amount of benefit it provides is not uniform and is insufficient to prove viable in the very low resource cases. Banyan is able to learn competitive representations consistently across languages, unsupervised and with very little data, providing a viable alternative for cheaply and efficient embeddings for low resource languages.
5.3 Efficiency
Other than its embedding matrix, Banyan only has composition and decomposition functions. Diagonalising these makes them easier to compute and more lightweigh than standard weight matrices, ($2U$ rather than $2U\times U$), achieving a further order of magnitude reduction in parameters compared with the already minimal Self-StrAE.
Secondly, by exploiting entangled tree structure the number of nodes grows at a significantly reduced rate with batch size compared with standard sentential trees (see Figure 4). This is because the number of reused constituent nodes also grows as batch size increases, and entangled trees capture the set of all constituents, which consequently does not grow as drastically. In practical terms, because entangled trees requires fewer nodes, and each node requires two distinct embeddings ($\bar{e}$ and $\underline{e}$) to be held for it, reducing the number of nodes required leads to radical improvements in memory efficiency. Put together, these changes mean that we can train Banyan very quickly as we can use large batches and its small number of parameters ensure quick convergence. On a single Nvidia A40 GPU with a batch size of 1024, Banyan pretrains from scratch in under 50 minutes, meaning that the total cost of pretraining a Banyan model sits at around 30 cents [^2] Cloud computing costs from: https://www.runpod.io/pricing. Free-tier Google Colab users can achieve similar results in about two hours with a smaller batch size. Inference can also be performed on CPU on typical laptops, because the model is so small. Combined with its data efficiency, we believe this provides a promising alternative for low resource languages and communities.
@@ 5.3 Efficiency
+ diag functions 44.15 $\pm$ 0.002 62.80 $\pm$ 0.002 48.30 $\pm$ 0.001 64.60 $\pm$ 0.002 60.30 $\pm$ 0.001 49.80 $\pm$ 0.002 55.14 $\pm$ 0.001 57.70 $\pm$ 0.001 55.23 $\pm$ 0.001
++ CE loss 51.20 $\pm$ 0.007 69.10 $\pm$ 0.002 63.20 $\pm$ 0.004 73.20 $\pm$ 0.002 66.60 $\pm$ 0.002 61.50 $\pm$ 0.002 55.50 $\pm$ 0.003 61.60 $\pm$ 0.002 62.70 $\pm$ 0.001
Table 6: Number of non-embedding parameters.

### v4

@@ Abstract
We present Banyan, a model that efficiently learns semantic representations by leveraging explicit hierarchical structure. While transformers excel at scale, they struggle in low-resource settings. Conversely recent structured models have shown promise as efficient learners, but lack performance. Banyan bridges this gap with two key innovations: an entangled hierarchical tree structure and diagonalised message passing, enabling it to outperform larger transformer models with just 14 non-embedding parameters. It excels in low-resource settings, offering a viable alternative for under-represented languages and highlighting its potential for efficient, interpretable NLP in resource-constrained environments.
@@ 1 Introduction
Banyan, achieves performance comparable to transformer-based baselines and represents a low-cost, viable alternative to transformers for producing representations in low-resource languages, as measured by semantic textual similarity (STS) tasks. By leveraging cognitively inspired inductive biases, our work enables semantic representation learning that rivals or surpasses large-scale pre-trained LLMs—using only 14 non-embedding parameters. Our model, Banyan, offers a new direction for efficient and effective semantic understanding in resource-constrained environments.  [^1] Code available at: github.com/exlab-research/Banyan
@@ 2 Background and Related Work
Recursive Neural Networks (RvNNs): Like regular recurrent neural networks (RNNs), RvNNs process data by repeatedly applying a function to update their state in sequence. However, instead of relying on temporal ordering (like the sequence of words in a sentence), RvNNs use hierarchical structures, often provided as input—most commonly as a binary tree—and can be applied either bottom-up (from leaves to root) or top-down (root to leaves). They were popularised in the deep learning era by Socher et al. (2011, 2013), inspiring many successor models that vary in how they define the recursive function, including Tree-LSTMs (Tai et al., 2015) and IORNN (Le & Zuidema, 2014).
@@
Semantic Representations of Text: Systems like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) use the distributional hypothesis (Harris, 1954) to model word semantics, which posits that words are defined by the context in which they appear. To learn representations, these models use a fixed context window and predict a missing word in a sequence. While initially effective, this approach is limited because representations for higher level objects (i.e. phrases, sentences etc.) are computed by simply averaging word embeddings. However, some notable follow on works attempted to improve upon this. Arora et al. (2017); Ethayarajh (2018) introduce a more sophisticated form of taking an average over word embeddings by using SVD, while Rücklé et al. (2018) use the power mean. These approaches yielded improvements, but relied on representations pre-trained at scale which limited their applicability to specialised domains or low resource languages. Wieting et al. (2021) attempt to refine the sentence representation through the use of paraphrase corpora, looking to increase alignment between language pairs, again improving performance, but requiring large scale parallel corpora. Finally, Pagliardini et al. (2017) realised that if the average of the embeddings was going to be used to create the sentence representation, it makes sense to optimise it directly. Consequently, they modified the pre-training objective in order to have the average predict a missing word from a sentence. At scale this proved tremendously effective. However, despite offering substantial improvements, all these methods require scale and more importantly do not tackle the central limitation of word embeddings - the inability to handle changes in meaning dependent on context.
On the other hand, transformers, through self-attention, are able to represent contextualised meanings. However, early encoder-only transformer models produced poor representations (Reimers & Gurevych, 2019), especially compared to the more sophisticated approaches based on word embeddings. This was largely due to the anisotropy issue (Godey et al., 2024), which required the development of techniques using contrastive fine-tuning (Gao et al., 2021) to finally remedy. These approaches eventually surpassed word embeddings, and have become the method of choice for producing semantic representations. However, they still rely heavily on scale for success, as contrastive refinement is a final fine-tuning step applied to pre-trained models rather than directly incorporated within pre-training.
Semantic Representation Learning through Structure: Transformer embeddings have become more successful than static word embeddings due to their ability to handle varying contextual influences. Unlike attention mechanisms in transformers, which route information based on token relationships, some approaches use explicit graphs or structures, such as dependency (Levy & Goldberg, 2014; Vashishth et al., 2019) or constituency parses (Pham et al., 2015), to determine the focus of context windows. These models have the potential to bridge the gap between the efficiency of word embeddings and the contextualisation offered by transformers. This is because the discrete structure provides an input specific routing order which dictates interactions between atoms and consequently determines their influence on higher level representations - allowing for more flexibility than simple averaging. Most related to our work, Opper et al. (2023b) introduce two models. StrAE, which use constituency parsers to learn sentence-level embeddings alongside word embeddings, and Self-StrAE, which learns its own structure using representations. This latter model, Self-StrAE, serves as the foundation for Banyan and is described next.
@@ 4 Banyan
Figure 2: Entangled trees: Example of disjoint trees being transformed into an entangled trees.
@@ 4.1 From trees to entangled trees
We wish to have composition embeddings amortise over all possible contexts, and simultaneously, all decompositions embeddings to resolve to the same thing. The representation of an entity “Lisa” should encapsulate everything she could possibly eat. Simultaneously, the average of everything she could eat we should get back to “Lisa”. Self-StrAE does not explicitly model this behaviour in its structure. Decomposition embeddings of the same entity only interact when we calculate the loss. On top of this, because the loss is taken over the batch, they are actually treated as false negatives to each other. Even though they are terms that ought not be pushed away, the objective ask them to be.
@@ 4.2 Simplified Message Passing
with sigmoid non-linearity ($\sigma$) applied to parameters both for numerical stability and to enforce a decayed memory over structure depth. Repeated application of the diagonal composition function will decay the influence of nodes further down in the tree, thereby respecting its compression order. During composition representations can increase in magnitude as they are the sum of the two children. During decomposition representations will, by necessity, reduce back down in magnitude towards the leaves. Further mimicking the information flow specified by the entangled trees. Finally, they restrict encoder (comp) and decoder (decomp) embeddings to remain in the same space. Which makes amortisation required for successful reconstruction, letting us switch objective to cross entropy over the vocabulary. We provide analysis to support this claim later in § 7.
These relatively simple changes have a pretty drastic effect, both in terms of performance (see experiments), as well as efficiency, with parameters now reduced by a factor of $U$ compared to the functions from Equations 1 and 2.
Table 1: Sentence level results for models pretrained on English. Higher is better. Results represent mean and standard deviation across four random initialisations. Spearman’s $\rho$ is * 100 following convention.
@@
Banyan 14.65 $\pm$ 2.90 63.23 $\pm$ 2.21 67.73 $\pm$ 0.3 48.53 $\pm$ 1.33
5 Experiments: English Evaluation
Goal: We wish to test whether Banyan can efficiently learn semantics. We start by evaluating on English, which is well resourced and has a wide array of test sets available with which we can measure the efficacy of our embeddings. This is crucial to establish, because when we turn to low resource languages later on, the amount of reliable evaluation sets will become limited. We want to make use of broad spectrum of tests available for English to reliably demonstrate embedding quality before moving forward.
5.1 Experimental Setup and Evaluation:
We want to evaluate how well Banyan learns effective semantic representations. Ideally we want to probe this at different levels of hierarchy, because it allows us to test whether structured models can do what they are supposed to i.e., seamlessly transfer semantic knowledge across different levels of hierarchy via composition. Our evaluation is unsupervised, both to directly probe the effect of the inductive bias, and for greater parity with what may be expected in a low-resource domain. It consists of three parts:
Correlation with human judgements: We compare the cosine similarity of embedding pairs produced by the model with human judgements of their semantic correspondence. On the word level, we use Simlex-999 (Hill et al., 2015) and WordSim-S/R (Agirre et al., 2009). All tasks measure semantics, but do so on differing axes. Simlex measures similarity at the exclusion of relatedness. Wordsim S measures similarity without penalising relatedness. And Wordsim R measures relatedness. On the sentence level, we use STS-12 through 16 (Agirre et al., 2012, 2013, 2014, 2015, 2016), the STS-B (Cer et al., 2017), SICK-R (Marelli et al., 2014) and SemRel (Ousidhoum et al., 2024) - which combined cover a wide array of semantic correspondence.
Retrieval: This is a cornerstone of Retrieval-Augmented Generation (RAG)-based systems and perhaps the most important use case for embedding models. We use two retrieval datasets from the BEIR suite (Thakur et al., 2021). Quora: evaluates success of matching questions to answers and capturing the response relation. Arguana: evaluates matching arguments to counter arguments, testing if our semantic space captures the notion of dialectical opposition.
Classification: We also include two test sets from the GLUE benchmark (Wang et al., 2019). Sentiment classification (SST-2) tests whether the representation space captures semantic polarity. Paraphrase detection (MRPC): tests whether our representation space capture semantic equivalence. While our other evaluation is applied to the embeddings zero-shot, for the classification tasks we train a GeLU MLP with a 512D hidden size, though we leave models frozen as a direct test of representation quality.
@@ 5.1 Experimental Setup and Evaluation:
Baselines: We compare against the Self-StrAE, GloVe (Pennington et al., 2014), Sent2Vec (Pagliardini et al., 2017) and a RoBERTa (Liu et al., 2019) in the medium configuration from (Turc et al., 2019; Opper et al., 2023a). Self-StrAE, the closest point of comparison to Banyan, indicates where the current performance level of structured representation learning lies. GloVe lets us compare to traditional static embeddings, and tests whether our model is learning anything more than just simple bag of word features. To obtain sentence embeddings, we report results using both the simple average of the word embeddings and the average with stopwords removed following (Reimers & Gurevych, 2019). The latter is generally stronger. An even more powerful variant is Sent2Vec, which directly optimises the averaged representation by using it to represent the context. This comparison measures the utility Banyan’s flexible and parametrised composition process, compared to an optimised average. Finally, for RoBERTa, we report results using both the standard model, and again after enhancing RoBERTa through an extra round of contrastive SimCSE training (Gao et al., 2021), as a further non-lexical baseline. Our pooling strategy is mean. To produce static embeddings from RoBERTa to use in lexical evaluation, we follow Bommasani et al. (2020) and average the contextualised representations of all occurrences of the word in the training set. The RoBERTa is intended as a stronger baseline. It has significantly more parameters than Banyan and can model meaning in context unlike GloVe and Sent2Vec.
@@
Hyperparameters and Pre-Training Details: For all models we set the embedding size to 256. For Self-StrAE we use the configuration of (Opper et al., 2023b) and set embeddings as square matrices (i.e., $K$=16 and $U$=16). For Banyan we set these values to $K$=128 and $U$=2, because the more independent channels we allowed the better the model seemed to perform. We refer the reader to Appendix A for ablations. We also note that because we can perform this reduction in channel size, the number of non-embedding parameters for Banyan drops to just 14, as these are directly proportional to $U$. The configuration for RoBERTa medium is 8 layers, 8 attention heads, 2048 dimensional feed forward layers, and relative positional embeddings. For Self-StrAE, Banyan and Sent2Vec we pre-train on a uniform subsample of English Wikipedia consisting of circa 10 million tokens. This represents the lower end of how many tokens might be available in a low resource setting, and allows us to test whether these methods are efficient. Meanwhile for GloVe and RoBERTa we pre-train on Wiki-103 (Merity et al., 2016), to ensure that they are not penalised by insufficient scale. Wiki-103 comprises circa 100 million tokens and therefore represents the very upper end of what might be available in a low resource setting. Further training details are in Appendix B.
5.2 Results:
@@
Results are shown in Tables 1, 2 and 3. On both the word level and sentence level Banyan does much better than Self-StrAE. We ablate the reasons for this in more detail later in the manuscript. Both models suffer on SimLex because they need to model both similarity and relatedness as the latter dictates merge (related concepts often compose together). However, the important thing to note is that the structured models effectively transfer the same performance from the word level to the sentence level. They can take advantage of composition, and transfer the meaning of the parts to understanding the meaning of the whole. The GloVe baseline is good on the word level, but does not generalise to the sentence level as well as the transformer, even when we give it stopword removal. Similarly, Sent2Vec is extremely strong on the word level, and while more effective than GloVe, neither approach seamlessly transfers semantic knowledge to different levels of complexity. Banyan can, and is able to achieve comparable or better performance than the SimCSE RoBERTa despite being much smaller and exposed to 10x less pre-training data. This means we have a structured model that remains efficient and cheap, and also effective at representation learning.
@@
Model Indonesian Arabic Telugu Marathi Hausa Afrikaans Spanish Amharic Hindi Score
XLM-R 46.7 31.6 46.3 55.7 4.1 56.2 68.9 57.3 52.7 46.61
Llama-3.1 (8B) 53.4 31.1 65.6 63.4 6.1 65.4 66.7 64.1 61.7 53.06
Mistral Nemo 50.7 20.1 57 52.3 1.8 58.3 66.2 53.2 55.8 46.16
@@
XLM-R (FT) 47.9 33.6 68.8 75.1 14.6 72.6 72.8 59.6 57.6 55.84
Banyan 41.90 42.28 71.58 66.38 49.68 79.35 60.88 66.40 61.63 60.01
$\pm$ 0.56 $\pm$ 1.57 $\pm$ 1.24 $\pm$ 0.84 $\pm$ 0.75 $\pm$ 0.65 $\pm$ 0.86 $\pm$ 0.90 $\pm$ 0.43 $\pm$ 0.35
6 Experiments: Multilingual Evaluation
@@
Goal: We’ve established that Banyan is an efficient learner. This implies potential use for languages that are not well covered by current NLP approaches. Here we test that.
6.1 Experimental Setup and Evaluation:
@@
Tasks: Learning semantic representations for low-resource languages remains an ongoing challenge. A core problem is not just the lack of training data, but also the lack of evaluation datasets. Recent work by Ousidhoum et al. (2024) has sought to address this issue, providing semantic relatedness test sets for several low resource Asian and African languages, evaluated by comparing embedding similarity to human judgements. This means we can Banyan’s ability on Indonesian, Arabic, Telugu, Marathi, Hausa, Afrikaans, Spanish, Amharic and Hindi—covering a broad spectrum of resource extent. For example, Spanish is generally well represented, while Hausa is considerably less so.
@@
Baselines: We select XLM-R (Conneau et al., 2019): a transformer encoder trained on 2TB of multilingual data. Llama 3.1 8B (Dubey et al., 2024): a decoder only LLM trained on 15 trillion tokens. Mistral Nemo 12B: a decoder only LLM designed with multi-lingual capacities in mind. In addition we also compare against two specialised embedding models from the sentence transformers range (Reimers & Gurevych, 2019): Mini-LM-L12-V2 and Paraphrase-XLM-R-Multilingual-V1. These are pre-trained encoders that have been finetuned on supervised datasets designed to produce high quality semantic representations. The baselines we select here are emblematic of the kind of models one might reach for in order to embed a corpus. For all baselines we use mean pooling following (Reimers & Gurevych, 2019). Finally, for parity we include an XLM-R baseline which is finetuned on the same corpora as Banyan.
@@
Banyan Pre-training and Hyperparameters: For Afrikaans, Spanish and Amharic we obtained corpora from Leipzig Corpora Collection (Goldhahn et al., 2012). For Amharic we utilised a MIT licenced pre-training set of 1 million sequences available at this link. Hausa data was sourced from Opus (Nygaard & Tiedemann, 2003). Each dataset consists of roughly 10 million tokens. We utilise a pre-trained BPE tokeniser for each language from the BPEMB Python package (Heinzerling & Strube, 2018). Banyan’s hyperparameters remain the same as before. For XLM-R we finetune for 100k steps with early stopping, using a linearly scheduled learning rate of 5e-5 with 10% of steps as warmup. XLM-R runs at batch size 128 across 4xA40 cards.
Table 5: Number of non-embedding parameters.
@@
+ diag functions 44.15 $\pm$ 0.002 62.80 $\pm$ 0.002 48.30 $\pm$ 0.001 64.60 $\pm$ 0.002 60.30 $\pm$ 0.001 49.80 $\pm$ 0.002 55.14 $\pm$ 0.001 57.70 $\pm$ 0.001 55.23 $\pm$ 0.01
++ CE loss 51.38 $\pm$ 0.15 69.60 $\pm$ 0.37 63.20 $\pm$ 0.28 73.08 $\pm$ 0.26 67.18 $\pm$ 0.56 61.90 $\pm$ 0.63 55.23 $\pm$ 0.13 61.88 $\pm$ 0.22 62.97 $\pm$ 0.03
6.2 Results
See Table 4. In Spanish, a well-resourced language with high coverage, the transformer baselines almost all outperform Banyan. However, as languages become lower resourced the picture changes, and Banyan outperforms or is comparable to the baselines. This even includes the multilingual XLM-R that has undergone supervised training. While finetuning XLM-R improves performance the amount of benefit it provides is not uniform and is insufficient to prove viable in the very low resource cases. Banyan is able to learn competitive representations consistently across languages, unsupervised and with very little data, providing a viable alternative for cheaply and efficient embeddings for low resource languages.
7 Improvements and Ablations
7.1 Efficiency
Other than its embedding matrix, Banyan only has composition and decomposition functions. Diagonalising these makes them easier to compute and more lightweight than standard weight matrices, ($2U$ rather than $2U\times U$), achieving a further order of magnitude reduction in parameters compared with the already minimal Self-StrAE. Table Table 5 shows the difference, including a comparison to the various baselines used throughout the paper. Despite its size Banyan remains competitive.
Secondly, by exploiting entangled tree structure the number of nodes grows at a significantly reduced rate with batch size compared with standard sentential trees (see Figure 4). This is because the number of reused constituent nodes also grows as batch size increases, and entangled trees capture the set of all constituents, which consequently does not grow as drastically. In practical terms, because entangled trees requires fewer nodes, and each node requires two distinct embeddings ($\bar{e}$ and $\underline{e}$) to be held for it, reducing the number of nodes required leads to radical improvements in memory efficiency. Put together, these changes mean that we can train Banyan very quickly as we can use large batches and its small number of parameters ensure quick convergence. On a single Nvidia A40 GPU with a batch size of 1024, Banyan trains from scratch in under 50 minutes, meaning that the total cost of pretraining a Banyan model sits at around 30 cents [^2] Cloud computing costs from: https://www.runpod.io/pricing. Free-tier Google Colab users can achieve similar results in about two hours with a smaller batch size. Inference can also be performed on CPU on typical laptops, because the model is so small. Combined with its data efficiency, we believe this provides a promising alternative for low resource languages and communities.